14. Lamar Corporation sells a single product for $12 per unit. Last year, the company's sales revenue was $255,000 and its net operating income was $16,000. If fixed expenses totaled $82,000 for the year, the break-even point in unit sales was:

A. 15,000

B. 9,900

C. 17,781

D. 12,000
Answer #1
Units sold last year 21250 =255000/12
Sales revenue 255000
Less: Fixed expenses 82000
Less: Net operating income 16000
Variable expenses 157000
Unit Variable expense 7.39 =157000/21250
Fixed expenses 82000
Divide by Unit contribution margin 4.61 =12-7.39
Break-even point in unit sales 17781
Option C 17,781 is correct